OpenStudy (fanduekisses):

ACT Question: If x < 0 and (x+3)(x+1)=8, then x^2= A. -5 B. 2 C. 4 D. 5 E. 25

geerky42 (geerky42):

Though it would takes times, if I were you, I will just have (x+3)(x+1)=8 to a^2+bx+c = 0 form, then I'd try to factor it first, if failed, then use quadratic formula

OpenStudy (fanduekisses):

I get (x+5)(x-1)

OpenStudy (fanduekisses):

do they just want me to use any of the two x's? that's what confuses me sometimes lol

OpenStudy (fanduekisses):

because the correct answer is E.25 so I guess I just needed to use 5?

OpenStudy (fanduekisses):

oops nvm I think I am more confused lol

geerky42 (geerky42):

(x+5)(x-1) = 0, so what is x?

OpenStudy (fanduekisses):

x=-5, x=1 ?

OpenStudy (fanduekisses):

ooooh they say x<0

geerky42 (geerky42):

Yes, now read question, you do notice what x have to be?

geerky42 (geerky42):

yeah, so x have to be -5 only

geerky42 (geerky42):

So x^2 = (-5)^2 = 25

OpenStudy (fanduekisses):

yeah ^_^ thanks <3

geerky42 (geerky42):

welcome!